Hamilton Tours offers small group tours of the local area. The cost of a tour is $119 per person and the variable costs are $46 per person. Fixed costs for the quarter are $191,698.

Required:

a. What is the quarterly break-even level for Hamilton Tours?

b. Based on the sales director's forecast for the following quarter and using the price and cost information on the tours given, the margin of safety percentage is 50 percent. How many tours does the sales director plan to sell next quarter?

Required A

: What is the quarterly break-even level for Hamilton Tours?

Break-even level = ??? tours

Required B

: Based on the sales director's forecast for the following quarter and using the price and cost information on the tours given, the margin of safety percentage is 50 percent. How many tours does the sales director plan to sell next quarter?

Number of tours = ???

Which answer should be in ???
Answer #1

Given, Selling price of tour = $119 per person

Variable cost = $46 per person

Fixed cost for the quarter = $191698

So, contribution margin = Price - VC = 119-46 = $73.

a). break-even quantity is at a point when fixed cost equals the contribution margin of the trips

So let number of break-even trip be n

So, 191698 = 73*n => n = 2626 trips

Break-even level = 2626 tours in a quarter.

b). Given margin of safety for quarter = 50%

Margin of safety = (Budgeted Sales - Break-even sales)*100/Budgeted sales

So, 50 = (Budgeted sales - 2626)*100/Budgeted sales

So, Budgeted sales = 2*2626 = 5252 tours.
Answer #2

or 2,626/50%= 5,252        100-50= 50%